OpenStudy (anonymous):

At what point will the graphs of the two equations intersect x+y=1 10x+5y=7 a.(2/5, 3/5) b (2/3, 1/3) c.(1/5, 4,5) d.( 1 3/4, -3,4)

OpenStudy (anonymous):

Its option a

OpenStudy (anonymous):

We need to solve two equation Are you familiar with solving equation ?

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

Let me explain it x + y = 1 If we multiply this equation by 5 then we get 5x + 5y = 5 getting this ?

OpenStudy (anonymous):

Now we need to subtract this equation from second equation 

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

correct so x = ( 2/5) plug this value into first equation which is x + y = 1 can you try for this ?